Npmjs Status Updates

503s on www

Oct 21, 23:13 UTC Resolved - Our primary cache server for the website failed. The site is supposed to fail over to the spare cache server automatically, but due to a bug this didn't happen. Instead, we manually failed over to the spare server. The website was totally unavailable for 15 minutes from 15:47 to 16:02 Pacific Time. Our apologies!

About 998 days ago - Wed 21 Oct 2015 23:13 (GMT)

Source: Npmjs - npm

Npmjs History

503s to European users

Oct 18, 16:11 UTC Resolved - Networking issues caused users in Europe to receive higher numbers of HTTP 503 responses (caused by network timeouts to servers in the US) from 15:31 to 16:02 GMT today. These issues are now resolved.

About 1001 days ago - Sun 18 Oct 2015 16:11 (GMT)

More info

Elevated 503 rates on WWW

Oct 16, 07:40 UTC Resolved - From 06:46 to 07:31 UTC the website was returning elevated rates of 503 errors. The issue was resolved by a configuration fix and the website is now operating correctly.

About 1004 days ago - Fri 16 Oct 2015 07:40 (GMT)

More info

Public replication endpoint under heavy load

Oct 2, 21:09 UTC Resolved - This migration is now complete and all is well.Oct 2, 20:15 UTC Monitoring - is again operational. If your DNS is resolving to, then you are communicating with the new host. We're continuing to monitor.Oct 2, 19:54 UTC Update - The public skimdb registry database is briefly offline while we cut over to the newer hardware.Oct 2, 18:00 UTC Identified - We have identified that our public replication endpoint at is under heavy load due to organic traffic growth. We are moving to bigger hardware in the next 24 hours, which will result in a brief interruption in replication but all followers should recover automatically with no need for intervention. We will update this incident once the migration is complete.

About 1017 days ago - Fri 2 Oct 2015 21:09 (GMT)

More info

Empty tarballs responses

Sep 16, 16:29 UTC Resolved - Cache invalidation is complete.Sep 16, 15:55 UTC Monitoring - Cache invalidation is ongoing and should be complete within 15 minutes.Sep 16, 15:25 UTC Identified - A bug in code we rolled out to our canary server at 10:52 UTC today, and which our automated checks haven't caught caused us to return empty tarball responses to some of our users. We rolled the deployment back at 14:29 UTC and are working on invalidating caches affected by this outage.

About 1033 days ago - Wed 16 Sep 2015 16:29 (GMT)

More info

Elevated 503 rates in Australia

Sep 14, 21:42 UTC Resolved - Network issues have been resolved. We continue to work on longer-term architectural changes to be more resilient to trans-pacific network problems.Sep 14, 02:19 UTC Identified - We are working with our CDN to address the issue with our Australian PoP. In the meantime, Australian users can temporarily bypass the local PoP and go directly to a US PoP by setting their hosts file to point to -- one of our US POP IPs. If you use this workaround, please remember to reset your hosts file once the incident is resolved, as PoP IPs change periodically.Sep 14, 00:19 UTC Investigating - We are seeing an elevated level of failures (503s and network timeouts) affecting users near our CDN PoP in Australia. We are actively investigating.

About 1035 days ago - Mon 14 Sep 2015 21:42 (GMT)

More info

Older packages temporarily unavailable

Aug 5, 18:50 UTC Resolved - A security upgrade released yesterday accidentally prevented the registry from serving ~200 very old packages whose package data was incomplete. These packages combined account for less than 0.02% of registry downloads, so their errors did not trigger our monitoring overnight. Following user reports, we corrected the error this morning. The packages were unavailable for a period of ~14 hours, from 5pm PT 2015-07-04 to 7am PT 2015-08-05. Even though the absolute numbers involved are small, the absence of these packages severely disrupted some users and we apologize. We are putting additional auditing into place around releases to ensure this kind of accident cannot occur in future.

About 1075 days ago - Wed 5 Aug 2015 18:50 (GMT)

More info

External replication endpoint is unavailable

Jul 31, 20:53 UTC Resolved - All views are now up to date. Please contact if you have any issues get your mirrors back in sync. We have initiated some engineering work to make recovery from this kind of hardware failure much faster in future.Jul 31, 16:49 UTC Update - The replication endpoint at is now back up for most use-cases. Some views are still being populated.Jul 31, 04:03 UTC Monitoring - Restoring data is going to take several hours. Our apologies.Jul 31, 00:51 UTC Identified - A hardware failure has taken our external replication endpoint at offline. Engineers are working to replace this endpoint.

About 1080 days ago - Fri 31 Jul 2015 20:53 (GMT)

More info